OpenStudy (anonymous):

HELP!! I've been struggling through this all night! Find the equation of the line that is parallel to the graph of 2x-7y=6 and passes through (-1,-6). Write the equation in standard form.

OpenStudy (anonymous):

first, get the equation in standard form y=mx + b 2x - 7y = 6

OpenStudy (anonymous):

I've got that. I keep coming up with the wrong answer. I have y=2/7x-6/7

OpenStudy (anonymous):

okay, the equation is now in standard form the parallel line has the same slope like the line we were given

OpenStudy (anonymous):

Right I know that.

OpenStudy (anonymous):

so what's the slope m= ?

OpenStudy (anonymous):

m=2/7

OpenStudy (anonymous):

I know how to find the slope. I am in college! It's not that area I am find an issue!

OpenStudy (anonymous):

I don't have all day! I need this now.

OpenStudy (anonymous):

now use this slope/point equation to get a second parallel equation y-y0 = m(x-x0) find an equation with it, and you're done.

OpenStudy (anonymous):

I did and it's not right.

OpenStudy (anonymous):

I've done that several hundred times already and it's wrong.

OpenStudy (anonymous):

(-1,-6) x: -1 y: -6 y-(-6) = m(x- (-1) )

OpenStudy (anonymous):

y+6 = m( x + 1) y + 6 = mx + m y + 6 = 2/7 x + 2/7

OpenStudy (anonymous):

y = 2/7 x + 2/7 - 6 y = 2/7 x + 2/7 - 42/7 y = 2/7 x -40/7

OpenStudy (anonymous):

Wrong.

OpenStudy (anonymous):

you think so? it has the correct slope we can check the point to see if that point works in the equation P=(-1, -6) y = 2/7 x - 40/7 y = 2/7 (-1) -40/7 y = -2/7 -40/7 y = -42/7 y = -6 correct answer.

OpenStudy (anonymous):

It still didn't come out wrong and now I am starting from scratch again with a whole new set of numbers. I guess I just won't use this site since I'm not getting any help.